Assume the inflation rate is 4% per year and the market interest rate is 5% above the inflation rate.


Determine (a) the number of constant-value dollars 5 years in the future that is equivalent to $30,000


now and (b) the number of future dollars that will be equivalent to $30,000 now.

1
Expert's answer
2021-12-03T12:54:02-0500

The market interest rate is the rate of interest prevailing which is being offered on cash deposits. 


Using the formula,we find out nominal rate of interest,

nr=rr+i+(rr×i)

=12%+5%+(12×5)%

=17.6% or 0.176 

  

(where nr=nominal rate,rr=real rate of interest,i=inflation rate)

Using the formula,we find out future value,

FV=PV(1+r)


F=2000/(1+0.176)



Hence,correct option is B.